Suppose that X, Y, and Z are jointly distributed random variables, that is, they are defined on the same sample space. Suppose that we also have the following. E(x)-4 E(Y) 2 E(Z)-7 Var (x) -28 Var(Y)-3 Var (Z) -44 Compute the values of the expressions below. E(Y 1) 5Z + 4x Var (4Y-3)
Answer #1

E(-Y - 1) = -E(Y) - 1 = - 2 - 1 = -3

E((5Z + 4X)/(-3)) = (5*E(Z) + 4*E(X))/(-3) = (5*(-7) + 4*4)/(-3) = 19/3

Var(4Y - 3) = 16*Var(Y) = 16*3 = 48

E(-2X2) = -2E(X2) = -2(Var(X) + (E(X))2) = -2(28 + 42) = -88